7-(15 points) A nucleus decays into three channels with the relative 0.60, respectively. The half-life of the nucleus is 180 into the three channels? probabilities 0.10,030, and seconds. What are the partial decay rates
Answer #1

The total decay constant is

k=\frac{\ln 2}{t_{1/2}}=\frac{\ln 2}{180}=0.00385\,\,s^{-1}

Let the decay rates of the 3 channels be k_1,k_2,k_3

then the total decay rate

k_1+k_2+k_3=k

and given their relative probabilities, the partial decay rates are:

k_1 =0.1 k =0.000385\,\,s^{-1}=3.85\times 10^{-4}\,\,s^{-1}

k_2 =0.3 k =0.001155\,\,s^{-1}=1.155\times 10^{-3}\,\,s^{-1}

k 0.6k = 0.00231 s-1-2.31 × 10-3 s-